The growth rate of the FCFE of XYZ, Inc. is expected to be 15% per year for the next three years, after this three-year period, the growth rate in FCFE is expected to be 8% per year, indefinitely. Its trailing 12-month free cash flow to equity is $2.2.  XYZ’s cost of equity is 12% and weighted average cost of capital is 10%.  What is the stock’s intrinsic value estimate according to the two-stage FCFE Model?

A.

$104.58

B.

$71.26

C.

$64.30

D.

$75.09

E.

$61.97

F.

$62.16
Answer #1

Correct Answer is option B
Intrinsic value of FCFE = 71.26
To calculate the Intrinsic value,
Enter the stroke in the financial calculator in cash flow mode-
C1 = 2.53 (2.2 + 2.2*15% = 2.53)
C2 = 2.90950 (2.53 + 2.53*15z5 = 2.90950)
C3 =3.34593(2.90950 + 2.90950*15%) + terminal value
C3 = 3.34593+ 90.34 =93.68593
I = 12
CPT -NPV = 71.26
Calculation of terminal value = C3 + g / Re - g
After 3 year FCFE is grow at 8% forever
Cost of equity is taking rather than cost of capital, because we have to calculate FCFE or FCFF
terminal value = 3.34593 + 3.34593*8% / 0.12- 0.08
Terminal value = 3.61360 /0.04
Terminal value = 90.34

I hope this clear your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.